OpenStudy (anonymous):

If the length of one base of a trapezoid is 5, the length of the other base is 2x + 7, and the length of the midsegment is 6x – 24, what is the value of x?

OpenStudy (jmark):

Given length of one base of a trapezoid is 5, the length of the other base is 2x + 7, and the length of the midsegment is 6x – 24 than as we know => 2(6x - 24) = 5 + 2x + 7 =>12x-48=5+2x+7 =>10x=5+7+48 =>10x=60 => x = 6 CBSE Sample Papers for Class 9 Science http://goo.gl/V5qKzx